. Find the center and radius of the circle whose equation is x2 + 2x + y2 – 4y + 1 = 0.
Answer by stanbon(75887)

Find the center and radius of the circle whose equation is
x2 + 2x + y2 – 4y + 1 = 0.
------
Complete the square on the x and on the y terms::
x^2 + 2x + 1 + y^2 - 4y + 4 = -1 + 1 + 4
-------------
Factor::
(x+1)^2 + (y-2)^2 = 4
-----
Center = (-1,2)
Radius = sqrt(4) = 2
----------------------
Cheers,
Stan H.
-------------